OpenStudy (anonymous):

Multiply. (x-10) (x-10)

OpenStudy (anonymous):

HINT: (a-b)(a-b)=(a-b)^2 = a^2-2ab+b^2

OpenStudy (anonymous):

does it help?

OpenStudy (anonymous):

use distributive proerty